OpenStudy (anonymous):

Find an equation of the curve that passes thro. the pt (0,1) and whose slope at (x,y) is xy I would like a FAIR hint please, not the answer.

OpenStudy (anonymous):

Differential equations for sure

OpenStudy (anonymous):

@Mertsj @Zarkon @dpaInc @jim_thompson5910 @lgbasallote @KingGeorge @AccessDenied @SmoothMath @eliassaab @asnaseer @.Sam. @kropot72 @eigenschmeigen @satellite73 @amistre64 @TuringTest @shivam_bhalla @experimentX @FoolForMath @Callisto @phi

OpenStudy (anonymous):

Is this a tricky version of y = mx + b?

OpenStudy (anonymous):

Because I could do: y = (xy)*x + b

sam (.sam.):

the gradient or \(\frac{dy}{dx}\) is 'xy' you need to find the equation by integrating both sides

OpenStudy (anonymous):

I understand that "xy" will be on one side, but what will be on the other?

OpenStudy (anonymous):

Ok so the (0,1) portion is telling me that this is an INITIAL VALUE PROBLEM.

sam (.sam.):

\[\frac{dy}{dx}=xy\] \[\frac{1}{y}\frac{dy}{dx}=x\] \[\large \int\limits \frac{1}{y}\frac{dy}{dx} \, dx\text{ = }\int\limits x \, dx\] Do the math and then sub (0,1) to find the constant

OpenStudy (anonymous):

So this is some function s.t. when we take the derivative of at some point (x,y) we get xy

sam (.sam.):

yes, so when you integrate, you'll get back the equation or function

OpenStudy (anonymous):

Got it. Thanks.

sam (.sam.):

welcome